This problem is not from a book. I've solved every other problem on my worksheet but cannot seem to get this one. Please help!!!
Given f(x) = 12x^2 - 53x +55, solve f(x)"< or=" 0.
Express your answer using interval notation. If you need to use -infinity, type -INFINITY. If you need to use positive infinity, type INFINITY. Use fractions as needed.
Please note: I didn't know how to type "less than or equal" in this kind of message, so I put "< or=."
Thanks!



Answer by edjones(8007)

Given f(x) = 12x^2 - 53x +55
12x^2 - 53x +55<=0
Factor: (3x-5)(4x-11)<=0 or use quadratic formula.
5/3<=x<=11/4
[5/3,11/4]
Ed